1. Determine the volume of 0.7M hydrochloric acid needed to fully neutralize 60 mL of 0.5M Ca(OH)2.
  2. If 50 mL of 1.2M H3PO4 is neutralized by 75 mL of NaOH, what is the concentration of the NaOH?
  3. What volume of 0.2M HNO3 acid would be needed to fully neutralize 1L of 0.4M KOH?
  4. 200 mL of 2M Sr(OH)2 fully neutralized 60 mL of H2SO4. What is the concentration of the acid?
1
Expert's answer
2022-01-28T01:38:09-0500

Solution:

Molarity (M) = Moles of solute / Liters of solution

Moles of solute = Molarity × Liters of solution


Determine the volume of 0.7 M hydrochloric acid needed to fully neutralize 60 mL of 0.5 M Ca(OH)2.

Balanced chemical equation:

Ca(OH)2 + 2HCl → CaCl2 + 2H2O

According to stoichiometry:

Moles of Ca(OH)2 = Moles of HCl / 2

Molarity of Ca(OH)2 × Volume of Ca(OH)2 = Molarity of HCl × Volume of HCl / 2

Volume of HCl = (2 × Molarity of Ca(OH)2 × Volume of Ca(OH)2) / (Molarity of HCl)

Volume of HCl = (2 × 0.5 M × 0.06 L) / (0.7 M) = 0.0857 L = 85.7 mL

The volume of hydrochloric acid (HCl) is 85.7 mL



If 50 mL of 1.2 M H3PO4 is neutralized by 75 mL of NaOH, what is the concentration of the NaOH?

Balanced chemical equation:

H3PO4 + 3NaOH → Na3PO4 + 3H2O

According to stoichiometry:

Moles of H3PO4 = Moles of NaOH / 3

Molarity of H3PO4 × Volume of H3PO4 = Molarity of NaOH × Volume of NaOH / 3

Molarity of NaOH = (3 × Molarity of H3PO4 × Volume of H3PO4) / (Volume of NaOH)

Molarity of NaOH = (3 × 1.2 M × 50 mL) / (75 mL) = 2.4 M

The concentration of NaOH is 2.4 M


What volume of 0.2 M HNO3 acid would be needed to fully neutralize 1 L of 0.4 M KOH?

Balanced chemical equation:

HNO3 + KOH → KNO3 + H2O

According to stoichiometry:

Moles of HNO3 = Moles of KOH

Molarity of HNO3 × Volume of HNO3 = Molarity of KOH × Volume of KOH

Volume of HNO3 = (Molarity of KOH × Volume of KOH) / (Molarity of HNO3)

Volume of HNO3 = (0.4 M × 1 L) / (0.2 M) = 2 L

The volume of HNO3 is 2 L


200 mL of 2 M Sr(OH)2 fully neutralized 60 mL of H2SO4. What is the concentration of the acid?

Balanced chemical equation:

H2SO4 + Sr(OH)2 → SrSO4 + 2H2O

According to stoichiometry:

Moles of H2SO4 = Moles of Sr(OH)2

Molarity of H2SO4 × Volume of H2SO4 = Molarity of Sr(OH)2 × Volume of Sr(OH)2

Molarity of H2SO4 = (Molarity of Sr(OH)2 × Volume of Sr(OH)2) / (Volume of H2SO4)

Molarity of H2SO4 = (2 M × 200 mL) / (60 mL) = 6.667 M = 6.7 M

The concentration of the acid (H2SO4) is 6.7 M